### v1

# The neutron decay anomaly, neutron stars and dark matter

Abstract
The discrepancies in different measurements of the lifetime of isolated neutrons could be resolved by considering an extra neutron decay channel into dark matter, with a branching ratio of the order of $O(1$%). Although the decay channel into a dark fermion $\chi$ plus visible matter has been already experimentally excluded, a dark decay with either a scalar or dark photon remains still a possibility. In particular, a model with a fermion mass $m_{\chi}\approx 1$ GeV and a scalar $m_{\phi}\approx O(\rm{MeV})$ could provide not only the required branching ratio to explain the anomaly but also a good dark matter (DM) candidate with the right thermal abundance today. Although the interaction DM-neutron will affect the formation of neutron stars, the combined effect of the dark matter self-interactions mediated by the light scalar and an effective repulsive interaction with the neutrons induced by the scalar-Higgs coupling would allow heavy enough neutron stars. The combined constraints from neutron lifetime, dark matter abundance, neutron star and Higgs physics, and Big Bang Nucleosynthesis, restrict the light scalar mass to the range $0.03752m_{e}<m_{\phi}<2m_{e}+$ 0.0375 MeV.
I Introduction
The Standard Model (SM) of particle physics provides a successful description of the fundamental particles and their interactions. Nevertheless, we know that this is still an incomplete picture, as it does not explain for example the nature of Dark Matter (DM), the origin of the asymmetry matter-antimatter, or that of light neutrino masses. Besides, there are still open questions regarding some particle measurements, like the lifetime of the neutron, one of the building blocks of baryonic matter. The isolated neutron mainly decays via $\beta$ processes into protons, electrons and anti-neutrinos:
$n\rightarrow p+e^{-}+\bar{\nu}_{e}\,.$ (1)
Two different methods have been used to measure this lifetime. The bottle or storage method Mampe1993 ; Serebrov2005 ; Serebrov2008 ; Pichlmaier2010 ; Steyerl2012 ; Arzumanov2015 counts the number of neutrons that have not decayed, with an average lifetime value $\tau_{n}^{\text{bottle}}=878.4\pm 0.6\,$s ParticleDataGroup2022 . On the other hand, the beam method Byrne1996 ; Yue2013 ; Sumi:2021svn counts protons from $\beta$ decay, with an average $\tau_{n}^{\text{beam}}=888.0\pm 2.0\,$s Sumi:2021svn . The results are highly dependent on the method, and this discrepancy of $9.6$ s or $4.0\,\sigma$ is known as the neutron decay anomaly. Although it is not excluded that the difference is due to (yet unknown) systematic errors Giacosa2020 , it might also be due to a new decay channel unaccounted for in beam experiments. For example, as proposed in Ref. Fornal2018 , the discrepancy could be explained with a new neutron decay channel into a DM fermion $\chi$:
$n\rightarrow\chi+{\rm visible}({\rm invisible})\,,$ (2)
where “visible” can be either a photon or a pair of $e^{-}e^{+}$, and “invisible” either a dark photon ($A^{\prime}$) or a dark scalar ($\phi$). Explaining the difference in the neutron lifetime would require a branching ratio ${\rm Br}(n\rightarrow\chi\,+\,{\rm anything})\approx 1\%$. The “visible” channel with a photon was soon experimentally excluded Tang:2018eln , and the one with a pair of $e^{-}e^{+}$ constrained for kinetic energies $37.5\,\rm{keV}\lesssim E_{e^{+}e^{-}}\lesssim 644\,\rm{keV}$ (95 % CL) UCNA:2018hup ; Klopf:2019afh . For the invisible dark decay channel, the proposal was questioned related to Neutron Stars (NS) physics tovns1 ; tovns2 ; tovns3 . The mixing between neutrons and dark fermions leads to the partial conversion of neutrons to DM inside the star, softening the equation of state (EoS) at high densities, and lowering the possible maximal mass of the NS below the largest observed mass $M_{NS}\sim 2M_{\odot}$ Riley2021 . However, as it was already noticed in tovns1 , DM repulsive self-interactions could avoid the problem. For example in Cline2018 the dark sector was assumed to be charged under an extra $U(1)$ symmetry, and extended including a dark photon $A^{\prime}$. The dark repulsive Coulomb-like interaction was enough to compensate for the softening of the EoS and to recover large enough NS masses; however different cosmological, astrophysical and particle physics constraints preclude the possibility of the dark fermion accounting for the whole DM abundance in this model. Another possibility was explored in Grinstein2019 where they considered the invisible decay channel with a singlet dark scalar. Through the mixing of the dark scalar with the Higgs, an effective repulsive DM-neutron interactions can be generated, which again would disfavour the production of DM inside the NS, and this model could also account for the right DM relic abundance. Another simpler possibility to recover massive enough NS and the dark matter relic abundance was proposed in Strumia:2021ybk ; Husain:2022brl , which considered the decay into 3 dark fermions.
In this work we focus on the model presented in Grinstein2019 with a light scalar mediator $\phi$, and their explicit particle model realisation including the coupling to the Higgs sector. They worked with $m_{\phi}\approx O(0.1)$ eV, which requires extending further the model to ensure the scalar decays before Big Bang Nucleosynthesis (BBN); otherwise it would contribute to the effective number of relativistic degrees of freedom at the time of BBN, which is severely constrained by cosmological observations Pitrou:2018cgg ; Planck2020 . On the other hand, a scalar mediator with a mass $m_{\phi}>2m_{e}$ could decay into a pair $e^{-}e^{+}$, avoiding BBN constraints. However, this will indeed reintroduce the neutron decay channel with a pair $e^{-}e^{+}$ in the final state
$n\rightarrow\chi+\phi^{*}\rightarrow\chi+e^{-}+e^{+}\,.$ (3)
Although this has been constrained to have a branching ratio less than 1% for $e^{-}e^{+}$ kinetic energies in the range $37.5\,\rm{keV}\lesssim E_{e^{+}e^{-}}\lesssim 644\,\rm{keV}$ at 95 % CL, still there is an unexplored region below and above this range, and we could have an invisible channel either with a dark light scalar with a mass $37.52m_{e}<m_{\phi}\lesssim 2m_{e}+$ 37.5 keV, or $m_{\phi}\gtrsim 2m_{e}+644$ keV. In addition, DM self-interactions mediated by a light scalar with $m_{\phi}\sim O(\rm{MeV})$ might help to solve the Small Scale Structure problems Tulin2013a ; Tulin2013 ; Kaplinghat2014 ; Kahlhoefer2017 . We stress that from the point of view of the scalar sector, its decay into SM fermions, besides being the simpler solution to the potential problems at BBN, may also open new venues of searching for possible experimental signatures either in neutron/baryon physics or Higgs physics.
We will thus revise the different constraints on model parameters for this proposal to explain the neutron decay anomaly with a DM fermion with $m_{\chi}\approx m_{n}$ and a singlet scalar mediator $m_{\phi}\approx O(\rm{MeV})$, such that (a) it is compatible with the observed DM abundance; (b) it can also alleviate the problems of the cold DM due to the self-interaction cross-section mediated by the scalar singlet; and (c) it allows for a repulsive (Yukawa potential) interaction between the neutron and the DM fermion and then massive enough NS. In section II we will first review the microscopic model proposed in Grinstein2019 in order to set the notation and mass ranges. In section III we compute the DM abundance from the standard freeze-out mechanism, and set the value of the coupling between the DM fermion and the light singlet scalar. Constraints from Higgs physics are revised in section IV, and those from NS physics in section V. Finally in section VI we summarize all the constraints and discuss the parameter space available. In Appendix A we just quote the expressions for the dark neutron decay.
II Particle Physics Model
The particle physics model proposed in Grinstein2019 is a modification of the one presented in Fornal2018 , and it requires two extra scalars and fermions: a heavy scalar $\Phi=(3,1)_{-1/3}$ with baryon number $B_{\Phi}$=-2/3, which mediated the neutron decay; a light scalar singlet $\phi$ with $B_{\phi}=0$ coupled to the Higgs, which generates the repulsive interaction; the DM Dirac fermion $\chi$, and an extra Dirac fermion $\tilde{\chi}$ heavier than the neutron, with $B_{\chi}=B_{\tilde{\chi}}=1$. At the quark level, the interactions are given by:
$\begin{split}\mathcal{L}=&(\lambda_{q}\,\epsilon^{ijk}\,\overline{u_{Li}^{c}}\,d_{Rj}\,\Phi_{k}+\lambda_{\chi}\,\Phi^{*i}\,\overline{\tilde{\chi}}\,d_{Ri}+\lambda_{l}\,\overline{Q_{Ri}^{c}}\,l_{L}\,\Phi^{*i}+\lambda_{Q}\,\epsilon^{ijk}\,\overline{Q_{Ri}^{c}}\,Q_{Lj}\,\Phi_{k}+\text{h.c.})\\ &-M_{\Phi}^{2}|\Phi|^{2}-m_{\tilde{\chi}}\,\overline{\tilde{\chi}}\,\tilde{\chi}-m_{\phi}^{2}|\phi|^{2}-m_{\chi}\,\overline{\chi}\,\chi\\ &+\lambda_{\phi}\,\overline{\tilde{\chi}}\,\chi\,\phi+\mu\,H^{\dagger}H\,\phi+g_{\chi}\,\overline{\chi}\,\chi\,\phi+g_{\phi H}\,\phi^{2}H^{\dagger}H+\text{h.c.}\,.\end{split}$ (4)
The trilinear interaction $\mu$ between the singlet and the Higgs gives rise to an effective $g_{n}\bar{n}n\phi$ vertex, with $g_{n}=\mu\sigma_{\pi n}/m_{h}^{2}$, where $m_{h}$ is the Higgs mass and $\sigma_{\pi n}=\sum_{q}\langle n|m_{q}\bar{q}q|n\rangle\approx 370\,$MeV Finkbeiner2008 , with the sum running over all quark flavours. This originates the repulsive (with $\mu<0$) effective potential between the neutron and the DM thorough the exchange of the light scalar $\phi$:
$V=\frac{g_{\chi}g_{n}}{4\pi}\frac{e^{-m_{\phi}r}}{r}\,.$ (5)
Except for the masses of $\chi$ and $\phi$, we assume that other particles in the model are always heavy enough to avoid constraints (if any) from accelerator physics or cosmology. In addition, as we will see later, the extra fermion $\tilde{\chi}$ will not affect the DM abundance when $m_{\tilde{\chi}}\gtrsim 50$ GeV. In order to allow for the neutron decay channel into $\phi+\chi$, but ensuring the stability of nuclei Pfutzner2018 , one must impose :
$937.993~{}\rm{MeV}<m_{\chi}+m_{\phi}<939.565~{}\rm{MeV}\,,$ (6)
where the upper limit is the neutron mass and the lower limit avoids the decay of ${}^{9}Be$. One must also impose $m_{\tilde{\chi}}>937.993$ MeV to avoid the decay of ${}^{9}Be$ through $n\rightarrow\tilde{\chi}\gamma$. Eq. (6) ensures the stability of the proton, and to prevent the possible decay of the dark fermion into protons, i.e, $\chi\rightarrow p^{+}+e^{-}+\phi$, one needs:
$|m_{\chi}-m_{\phi}|<938.783~{}\rm{MeV}\,.$ (7)
Instead of working with a light singlet scalar $m_{\phi}\lesssim 2m_{e}$, which would require further extending the model to ensure their decay before BBN, we will work with $m_{\phi}\approx O(\rm{MeV})$, excluding the values constrained by the limits on the induced decay $n\rightarrow\chi+e^{-}+e^{+}$ set in UCNA:2018hup ; Klopf:2019afh . The DM fermion mass is then $m_{\chi}\sim O(1\,\rm{GeV})$, within the range given in Eq. (6). The condition in Eq. (7) is always fulfilled, and the fermion $\chi$ is stable and a good candidate for DM.
About the couplings, neutron dark decay rate will depend on the combination of couplings $\lambda_{\phi}\lambda_{q}\lambda_{\chi}$ Fornal2018 (see Appendix A), and requiring the branching ratio to be less than 1% only gives a mild constraint $10^{-5}\lesssim\lambda_{\phi}\lambda_{q}\lambda_{\chi}\lesssim\times 10^{-2}$ with $M_{\phi}\simeq m_{\tilde{\chi}}\simeq 1\,\rm{TeV}$. The couplings in the last line in (4) will be set in the following sections demanding the model to be consistent with the DM abundance, Higgs and NS physics.
III Dark matter abundance
The DM abundance can be derived from the standard freeze-out mechanism, integrating the Boltzmann equation Steigman2012
$\frac{dn_{\chi}}{dt}=-3Hn_{\chi}-<\sigma v>\left[n_{\chi}^{2}-(n_{\chi}^{eq})^{2}\right]\,,$ (8)
where $n_{\chi}^{\text{eq}}$ is the number density in equilibrium, $H$ the Hubble parameter and $<\sigma v>$ the thermal averaged cross section, usually parameterised as
$<\sigma v>=\frac{1}{m_{\chi}^{2}}\left(\alpha_{s}+\alpha_{p}\frac{T}{m_{\chi}}\right)\,,$ (9)
where $\alpha_{s}$, $\alpha_{p}$ are dimensionless parameters associated with s-, p-waves annihilation channels respectively. For the model given in Eq. (4), DM annihilation cross-section into light scalars is given by Kouvaris2015
$\sigma_{\bar{\chi}\chi\rightarrow\phi\phi}v_{\text{rel}}\simeq\frac{3g_{\chi}^{4}}{32\pi m_{\chi}^{2}}\cdot\frac{T}{m_{\chi}}\,,$ (10)
and therefore $\alpha_{p}=3g_{\chi}^{4}/(32\pi m_{\chi}^{4})$. The observational value of the DM density parameter $\Omega_{DM}$ Planck2020 ,
$\Omega_{DM}h^{2}=0.120\pm 0.001\,,$ (11)
where $h$ is the reduced Hubble parameter, is reproduced with $\alpha_{p}=1.78\cdot 10^{-7}$, which for $m_{\chi}\simeq 1$ GeV implies $g_{\chi}\simeq 0.05$. We notice that the freeze-out mechanism and Eq. (8) assume implicitly that both the dark matter and the annihilation products are initially in thermal equilibrium. For the singlet scalar, this can be achieved before electroweak symmetry breaking through the scalar-Higgs interaction with a coupling value $g_{\phi H}\gtrsim 10^{-7}$ Enqvist2014 . The DM fermion is then kept in equilibrium with the singlet until it decouples at a temperature $T_{D}\approx 20m_{\chi}\simeq 20$ GeV.
We have also computed the relic abundance with micrOMEGAS Belanger2010 ; Alguero:2023zol , including the effect of the heavier fermion $\tilde{\chi}$. The coloured scalar $\Phi$ will not play any role as far as it is heavy enough, and we have set $M_{\phi}=1$ TeV. Other model parameters are taken as: $\lambda_{\phi}=0.04$, $g_{\phi H}=5\times 10^{-3}$, and $\mu=40$ GeV. The results for different values of $m_{\tilde{\chi}}$ are shown in Fig. (1). We obtain the right abundance with $g_{\chi}\simeq 0.053$, same than the one from our Boltzmann calculation. As it can be seen on the LHS in Fig. 1, this value is practically independent of the value of $m_{\tilde{\chi}}$, and for $m_{\tilde{\chi}}\geqslant 50\,$GeV, the extra fermion $\tilde{\chi}$ does not have any impact on the relic abundance calculation. On the RHS plot in Fig. 1 we have fixed $m_{\tilde{\chi}}=100$ GeV and vary the coupling $\lambda_{\phi}$ among the dark fermions $\chi$, $\tilde{\chi}$ and the singlet scalar. The DM density parameter shows some dependence on the value of the coupling $\lambda_{\phi}$ but again this disappears for values $\lambda_{\phi}\lesssim 0.1$.
Figure 1: Results obtained with micrOMEGAS for the relationship between the coupling $g_{\chi}$ and the DM relic abundance, (a) for different values of $m_{\tilde{\chi}}$ and $\lambda_{\phi}=0.04$ (LHS plot); (b) varying the value of $\lambda_{\phi}$ for $m_{\tilde{\chi}}=100$ GeV (RHS plot). See the text for other model parameters. The vertical dashed line indicates the value $g_{\chi}\simeq 0.053$ required to get today’s DM abundance.
IV Scalar potential
We now check constraints on the parameters of the scalar potential from electroweak symmetry breaking and the limit on the invisible Higgs decays. This was done in Ref. Kouvaris2015 , and we follow here their approach for the parameter values we are interested in.
The general potential between a singlet real scalar $S$ and a doublet $\Sigma$ is given by:
$V(S,\Sigma)=m_{\Sigma}^{2}|\Sigma|^{2}+m_{S}^{2}S^{2}3+\lambda|\Sigma|^{4}+\frac{\lambda_{S}}{4}S^{4}+\frac{\lambda_{S\Sigma}}{2}S^{2}|\Sigma^{2}|+\frac{\mu_{3}}{3}S^{3}+\mu_{1}S|\Sigma^{2}|\,,$ (12)
where $m_{\Sigma}$, $m_{S}$ are mass parameters, $\mu_{1}$, $\mu_{3}$ trilinear couplings, and $\lambda$, $\lambda_{S}$, $\lambda_{S\Sigma}$ dimensionless quartic couplings. The physical states $h$ and $\phi$ are obtained through the mixing of the neutral component of the doublet $\sigma^{0}$ and the singlet $s$:
$\sigma^{0}$ $=$ $\cos\theta h+\sin\theta\phi\,,$ (13)
$s$ $=$ $\cos\theta\phi-\sin\theta h\,,$ (14)
with a mixing angle after electroweak symmetry breaking given by:
$\sin\theta\simeq\frac{v_{EW}}{m_{h}^{2}}(\mu_{1}+\lambda_{S\Sigma}w)\,,$ (15)
where $m_{h}=125$ GeV is the Higgs mass , $v_{EW}=246.22$ GeV its vacuum expectation value (vev), and $w$ the singlet one. With $m_{\chi}=g_{\chi}w\simeq 1$ GeV, one has $w\lesssim 19$ GeV, i.e., $w\ll v_{EW}$.
Given that $m_{\phi}\ll m_{\chi}\simeq 1$GeV, the Higgs can now decay into light scalars and dark fermions with rates [^1] We keep the notation $g_{\chi}$ for the coupling between the physical scalar $\phi$ and the fermion $\chi$, and therefore the coupling between the physical Higgs $h$ and $\chi$ is given by $g_{\chi}\tan\theta$. :
$\Gamma(h\rightarrow\phi\phi)$ $=$ $\frac{\lambda_{S\Sigma}^{2}v_{EW}^{2}}{8\pi m_{h}}\left(1-\frac{4m_{\phi}^{2}}{m_{h}^{2}}\right)^{1/2}\,,$ (16)
$\Gamma(h\rightarrow\chi\chi)$ $=$ $\frac{g_{\chi}^{2}\tan\theta^{2}m_{h}}{8\pi}\left(1-\frac{4m_{\chi}^{2}}{m_{h}^{2}}\right)^{3/2}\,.$ (17)
The invisible branching ratio must be less than 12% AtlasHiggs , and using the SM Higgs decay rate $\Gamma_{h}=4.1$ MeV, the decay into scalars sets an upper limit $\lambda_{S\Sigma}\lesssim 5\times 10^{-3}$. Similarly from Eq. (17) using $g_{\chi}\simeq 0.053$ we obtain the upper limit on the mixing angle $\tan\theta\lesssim 0.1982$ ($\sin\theta\lesssim 0.1944$). On the other hand, the decay rate of $\phi$ into a pair of $e^{+}e^{-}$ is given by:
$\Gamma_{\phi}=\frac{h_{e}^{2}\sin\theta^{2}}{8\pi}m_{\phi}\left(1-\frac{4m_{e}^{2}}{m_{\phi}^{2}}\right)^{3/2}\,,$ (18)
where $h_{e}=2.9\times 10^{-6}$ is the SM electron Yukawa coupling. To be safe from BBN constraints Pitrou:2018cgg ; Planck2020 , we need $\tau_{\phi}\lesssim 1$ s, which only requires a mixing angle $\sin\theta\gtrsim 10^{-6}$ Kouvaris2015 ; Kaplinghat2014 ; Lin2012 ; ArkaniHamed2009 ; Tulin2013 .
Given that $w<v_{EW}$ and $\lambda_{S\Sigma}\ll 1$, the singlet practically does not contribute to the Higgs mass, $m_{h}^{2}\simeq 2\lambda v_{EW}^{2}$, and $\lambda\simeq\lambda_{SM}\simeq 0.13$. For the trilinear coupling $\mu$ among physical states we have:
$\mu\simeq 4(\mu_{1}+\lambda_{S\Sigma}w)\simeq 4\frac{m_{h}^{2}}{v_{EW}}\sin\theta\,,$ (19)
where we have used Eq. (15). For the repulsive interaction among neutrons and DM fermions we also need $\mu<0$, i.e. $\mu_{1}<-\lambda_{S\Sigma}w$. Therefore, from the limits on the mixing angle, the absolute value of the trilinear coupling must lie in the range:
$2.5\times 10^{-4}~{}\rm{GeV}\lesssim|\mu|\lesssim 49.4~{}\rm{GeV}\,.$ (20)
V Neutron Stars and DM
For a stellar compact object like a NS, the Mass-Radius (MR) relation can be obtained by integrating Einstein equations for a static, spherically symmetric configuration, the Tolman-Oppenheimer-Volkoff (TOV) equations tov1 ; tov2 .
$\frac{dP(r)}{dr}$ $=$ $-\frac{G(\rho(r)+P(r))}{r^{2}}\cdot\frac{M(r)+4\pi P(r)r^{3}}{1-2GM(r)/r}\,,$ (21)
$\frac{dM(r)}{dr}$ $=$ $4\pi r^{2}\rho(r)\,,$ (22)
where $G$ is the Newton constant, $P$ the pressure, $\rho$ the energy density, $M$ the stellar mass and $r$ the radius. TOV Eqs. are integrated for a given initial pressure (or number density) at the center of the object, and a given equation of state, i.e., the relation between $P$ and $\rho$, until we reach the surface pressure $P(R)=0$, at which point we also get the total mass and the maximal radius $R$.
Astrophysical observations in principle can be used to put constraints on the EoS of the NS, excluding those that are not able to reach the $2M_{\odot}$ limit, the largest observed NS mass MassNS . Adding the new neutron decay channel, one has to take into account the conversion of the neutrons into the DM fermion inside the NS. And typically, the addition of the DM particle effectively modifies the EoS inside the stellar object such that the maximum NS mass falls well below the observational limit. One possible solution is to allow for repulsive DM interactions Cline2018 ; Grinstein2019 , which disfavours the conversion.
There is a variety of EoS for nuclear matter in the literature Pieper:2001ap ; Wiringa:2002ja ; Gandolfi:2011xu ; Hebeler:2013nza , and in order to check how the DM hypothesis modifies the NS mass-ratio relation we have chosen to work with a simple parametrization of the internal energy per neutron:
$E_{n}(x)=ax^{\alpha}+bx^{\beta}\,.$ (23)
where $x=n_{n}/n_{0}$, $n_{n}$ the nucleon density and $n_{0}=0.16~{}{fm}^{-3}$ the saturation density. Different sets of parameters correspond to different choices of the EoS, and we work with those given in Table 1.
Model Label $a$ (MeV) $\alpha$ $b$ (MeV) $\beta$
none QMC1 12.7 0.49 1.78 2.26
$V_{2\pi}^{PW}+V_{\mu=150}^{R}$ QMC2 12.7 0.48 3.45 2.12
$V_{2\pi}^{PW}+V_{\mu=300}^{R}$ QMC3 12.8 0.488 3.19 2.20
$V_{3\pi}+V_{R}$ QMC4 13.0 0.49 3.21 2.47
$V_{2\pi}^{PW}+V_{\mu=150}^{R}$ QMC5 12.6 0.475 5.16 2.12
$V_{3\pi}+V_{R}$ QMC6 13.0 0.50 4.71 2.49
UIX UIX 13.4 0.514 5.62 2.436
Table 1: EoS given by the QMC predictions of two-body and three-body interactions. The second column shows the name used in this work to identify the different equations.
We have first integrated the TOV Eqs. (21)-(22) without DM, for the set of EoS given in Table (1). The mass-radius curves are plotted on the LHS of Fig. (2). We have also included observational limits from NICER collab. Miller2019 ; Riley2019 ; Miller2021 ; Riley2021 , and the not yet confirmed limit from LIGO collab. GW190814 Abbott2020 . These examples serve us a guide to understand the physics when including DM. For our case of study, only the EoS UIX and QMC6 are able to reproduce NS with $2M_{\odot}$ with a radius within the observable range. We focus therefore on these cases, adding DM with self-interactions, and repulsive $n-\chi$ interactions. The total energy density of the NS is now given by:
$\rho_{T}(n_{n},n_{\chi})=\rho_{n}(n_{n})+\rho_{\chi}(n_{\chi})+\frac{n_{n}n_{\chi}}{2z_{n\chi}^{2}}\,,$ (24)
where $n_{\chi}$ is the dark fermion number density, the contribution from the repulsive potential is given by the last term, and $\rho_{\chi}$ is the dark fermion energy density
$\rho_{\chi}=\frac{m_{\chi}^{4}}{8\pi^{2}}\left(x\sqrt{1+x^{2}}(1+2x^{2})-\ln(x+\sqrt{1+x^{2}})\right)+\frac{n_{\chi}^{2}}{2z_{\chi}^{2}}\,,$ (25)
where $x=(3\pi^{2}n_{\chi})^{1/3}/m_{\chi}$, and the last term in Eq. (25) is the contribution due to DM self-interactions.
Object Mass ($M_{\odot}$) Radius (km)
Pulsar PSR J0030+0451 $1.44^{+0.15}_{-0.14}$ $13.02^{+1.24}_{-1.06}$ Miller2019
Pulsar PSR J0740+6620 $2.07^{+0.07}_{-0.07}$ $12.39^{+1.30}_{-0.98}$ Riley2021
$2.08^{+0.09}_{-0.09}$ $13.71^{+2.61}_{-1.50}$ Miller2021
GW190814 Abbott2020 $2.59^{+0.08}_{-0.09}$ $14.77-14.87$ KanakisPegios2021
$11.35-13.67$ Miao2021
Table 2: NS mass and radius observations with NICER (pulsars) and LIGO (GWs). It is not completely confirmed if the compact object in event GW190814 is the heaviest NS or a small black hole.
Figure 2: LHS: Mass-radius relation of NS described by the family of EoS seen in table 1, without DM. Although the differences among parameters of the same type are rather small, the resulting values for each EoS have non-negligible changes. The coloured contours represent the observational values on masses and radii from NICER and LIGO. RHS: Mass-radius relation including DM repulsive and self-interactions, as indicated by the value of $z_{n\chi}$ and $z_{\chi}$ in the plot; $g_{n\chi}=0$ means that no repulsive interaction is included. We only considered as examples the EoS QMC6 and UIX.
Figure 3: Parameter space available in the plane $z_{\chi}-z_{n\chi}$ when including DM. Grey areas labeled “UIX” are excluded when using the UIX EoS because they do not give rise to 2 solar masses NS with a large enough radius with either $R>11.4$ km (dark grey) or $R>12.2$ km (light grey). Similarly for the violet shaded region for the QMC6 EoS. The brown shaded region on the bottom is excluded because the large repulsive interaction will give rise to a “DM” star instead of a NS. This bound is independent of the particular EoS used for nuclear matter. Vertical solid lines are the values of $z_{\chi}$ for a given value of $m_{\phi}$ as indicated in the plot, and $g_{\chi}\simeq 0.053$ as required to have the right DM abundance obtained in section III. On the left, the narrow band between vertical red dashed lines labeled $m_{\phi}\sim 1$ MeV is the range not excluded by PERKEOII experiment. The left hand corner solid (dashed) blue line is the lower limit on $z_{n\chi}$ from Higgs physics, with Br(Higgs $\rightarrow$ invisible)=12 % (6 %). The white region above this line is the parameter space available fulfilling all constraints considered. See the text for details.
In term of masses and couplings at the particle physics level given in Eq. (4), the variables $z_{i}$ are given by:
$z_{n\chi}=\frac{m_{\phi}}{\sqrt{g_{\chi}g_{n}}}\,,\;\;\;z_{\chi}=\frac{m_{\phi}}{g_{\chi}}\,,$ (26)
The presence of DM, even with self-interactions, tends to soften the effective EoS of the neutron star, given rise to smaller and less massive NS Cline2018 . This effect can be compensated by the effective repulsive interaction parametrised by $z_{n\chi}$. This can be seen in the example plotted on the RHS in Fig. (2), for the EoSs UIX and QMC6. We include as solid lines the case without DM for comparison, one without repulsive interactions (obtained by setting $g_{n\chi}=0$), and including the latter with $z_{n\chi}=2z_{\chi}=100$ MeV. However, a too large repulsive potential will expel the neutrons, and produce instead a “DM” star, i.e., a configuration with $n_{\chi}\gg n_{n}$. Therefore, independently of the EoS of nuclear matter, this imposes a lower bound on $z_{n\chi}\gtrsim z_{\chi}/\sqrt{2}$, such that when $z_{\chi}\simeq\sqrt{2}z_{n\chi}$ we reach the same equilibrium configuration for NS than in the absence of DM. Varying $z_{\chi}$, $z_{n_{\chi}}$ we then obtain the parameter space available in order to reproduce NS with 2 solar masses and a radius larger than 11.4 km Riley2021 (12.2 km Miller2021 ), shown in Fig. (3). We stress that although the upper limit on the repulsive parameter $z_{n\chi}$ depends on the EoS used for nuclear matter, and may disappear or be relaxed for other parametrizations of the EoS, the lower limit is universal. We also notice that the larger the value of the self-interacting parameter $z_{\chi}$ the more difficult to get massive enough neutron stars due to the softening of the EoS, and although this could be compensated by increasing the repulsive parameter $z_{n\chi}$, this is not possible beyond some value $z_{\chi}\approx 110$ MeV because the repulsive interaction turns the star into a DM one. Therefore, once the value of $g_{\chi}$ is set in order to obtain the right DM abundance, this would always translate into an upper bound for the light scalar mediator mass. Using the value of the coupling $g_{\chi}\simeq 0.053$ obtained in section III, we can set the value of $z_{\chi}$ for different $m_{\phi}$ masses. Those are indicated by the vertical dashed lines in the figure. The narrow vertical band between the first two red dashed lines to the left corresponds to the range $2m_{e}<m_{\phi}<2m_{e}+0.0375\,\rm{MeV}$, not excluded yet by the experiments Klopf:2019afh . Finally, the solid blue line at the left upper corner indicates the lower limit on the parameter $z_{n\chi}$ set by the upper limit on the trilinear coupling $\mu$, obtained in section IV from the current limit on the branching ratio of the invisible Higgs decay, such that:
$126.4\lesssim\frac{z_{n\chi}}{m_{\phi}}\lesssim 5.6\times 10^{4}\,.$ (27)
For comparison, we also include with a dashed blue line how this limit would be affected with a more restrictive limit say Br(Higgs $\rightarrow$ invisible)=6 %. This would reduce the maximum allowed value for the mixing angle Higgs-singlet, and then the trilinear parameter $\mu$, which increases the slope of the limit in the $z_{n\chi}-z_{\chi}$ plane. Combining all constraints, we are left with the values of the parameters in the white area. In the plot we have chosen not to exclude values of $m_{\phi}<2m_{e}$, but we remark that to access this region we need to extend the model such that the light scalar decays before BBN.
VI Summary and discussion
Let us summarize the values of masses and couplings for the model considered, from different physical considerations:
  • Neutron decay anomaly: The neutron decay anomaly could be explained with a new neutron decay channel into a dark fermion and scalar. An explicit model was given in Grinstein2019 , Eq. (4). Masses must be in the range given in Eqs. (6) and (7), to allow neutron decay but ensuring the stability of nuclei and the proton. We have chosen to work with a dark fermion mass $m_{\chi}\simeq 1~{}\rm{GeV}$, and a light scalar $m_{\phi}\sim O(\rm{MeV})$, but larger than $2m_{e}$. This allows the scalar to decay into a pair electron-positron, with a small enough lifetime to avoid problems at BBN. This will reintroduce the neutron decay channel $n\rightarrow\chi+e^{-}+e^{+}$, excluded so far only for certain values of the kinetic energy $E_{e^{-}e^{+}}$. We can then adjust the value of $m_{\phi}$ in order to avoid those limits. Other particles in the model, like $\tilde{\chi}$ and $\Phi$ are assumed to be heavy enough, say O(1 TeV).
  • Dark matter abundance: We have checked semi-analytically and with micrOMEGAS that a dark fermion with a mass $m_{\chi}\simeq 1$ GeV is a good candidate for dark matter, and reproduces the present dark matter abundance with a coupling value $g_{\chi}\simeq 0.053$.
  • Higgs physics: With the inclusion of a light singlet scalar, the standard model Higgs can now decay into both light scalars or dark fermions. Demanding that the invisible branching ratio be less than 12% sets an upper bound on the singlet-Higgs mixing angle, while a lower bound can be set to ensure the decay of the light scalar before BBN: $10^{-6}\lesssim|\sin\theta|\lesssim 0.1944$. This translates into the allowed range for the trilinear interaction given in Eq. (20).
  • Neutron Stars: The model in Grinstein2019 gives rise to self-interactions for the dark fermion, through the exchange of the light scalar, and an effective repulsive interaction neutron-dark fermion through the Higgs portal. Both interactions will affect the NS mass-radius relation, derived by integrating the TOV Eqs. Having fixed $m_{\chi}$, $m_{\phi}$ and $g_{\chi}$, the parameter for the self-interactions is given by $z_{\chi}\simeq 19m_{\phi}~{}\rm{MeV}$, which translates into a lower limit for the repulsive parameter $z_{n\chi}^{min}\gtrsim 13.3m_{\phi}~{}\rm{MeV}$, independently of the EoS. However, Higgs physics constraints set the range of possible trilinear singlet-Higgs interactions, and therefore a more restrictive lower limit on $z_{n\chi}/m_{\phi}\gtrsim 126.4$.
All constraints together, for the EoSs we have considered, demanding to have NS with $M_{NS}\simeq 2M_{\odot}$ and $R>12.2$ km Miller2021 just leaves a very narrow range of possible values for the light scalar mass $2m_{e}<m_{\phi}<2m_{e}+0.0375\,\rm{MeV}$ (see Fig. (3)). On the other hand, only requiring to have massive enough NS with $R>11.4$ km, would open the parameter space for the UIX (QMC6) EoS for slightly larger masses $m_{\phi}\simeq 2-2.77$ MeV. In any case, NS physics alone sets an upper limit on the light scalar $m_{\phi}\lesssim 6$ MeV, which gets further reduced when combined with the constraints on the mixing singlet-Higgs.
The lower limit on $m_{\phi}$ comes from the minimal requirement to ensure the decay of $\phi$ before BBN into a pair $e^{+}e^{-}$. Smaller values for $m_{\phi}$ might be possible but require extending the model as in the original proposal Fornal2018 . Nevertheless, the model with the singlet decaying into SM particles is still viable.
Appendix A Neutron decay
For completeness, we summarize here the expressions of the neutron decay rate for the different dark channels.
(a) Decay into scalar and fermion: Starting from the lagrangian in Eq. (4), and integrating out the heavy states, neutron effective interactions with a dark scalar $\phi$ and fermion $\chi$ are given by the effective lagrangian
${\cal L}_{n\rightarrow\chi\phi}^{eff}=\frac{\lambda_{\phi}\epsilon}{m_{n}-m_{\tilde{\chi}}}\bar{\chi}~{}n~{}\phi^{*}\,,$ (28)
with $\epsilon=\frac{\beta\lambda_{q}\lambda_{\chi}}{M_{\Phi}^{2}}$ and $\beta=0.0144(3)~{}\rm{GeV}^{3}$ from lattice calculations. The decay rate is then;
$\Gamma_{n\rightarrow\chi\phi}=\frac{\lambda_{\phi}^{2}\epsilon^{2}}{16\pi}\sqrt{f(x,y)}\frac{m_{n}}{(m_{n}-m_{\tilde{\chi}})^{2}}\,,$ (29)
where $x=m_{\chi}/m_{n}$, $y=m_{\phi}/m_{n}$, and:
$\sqrt{f(x,y)}$ $=$ $((1-x)^{2}-y^{2})^{1/2}((1+x)^{2}-y^{2})^{3/2}\,,$ (30)
Working in the limit $m_{\tilde{\chi}}\gg m_{n}$, we can write:
$\Gamma_{n\rightarrow\chi\phi}=\frac{\kappa_{\phi}^{2}}{16\pi}\sqrt{f(x,y)}m_{n}\,,$ (32)
with
$\kappa_{\phi}\simeq\lambda_{\phi}\lambda_{q}\lambda_{\chi}\cdot\frac{\beta}{m_{\tilde{\chi}}M_{\Phi}^{2}}\simeq 1.44\times 10^{-11}\left(\frac{\rm{TeV}^{3}}{m_{\tilde{\chi}}M_{\Phi}^{2}}\right)\lambda_{\phi}\lambda_{q}\lambda_{\chi}\,.$ (33)
And for example for $m_{\phi}=2me+0.0375$ MeV, this gives the branching ratios for the upper and lower limits on $m_{\chi}$ from Eq. (6):
$Br(n\rightarrow\chi\phi)$ $\simeq$ $16.3~{}(\lambda_{\phi}\lambda_{q}\lambda_{\chi})^{2}\left(\frac{\rm{TeV}^{3}}{m_{\tilde{\chi}}M_{\Phi}^{2}}\right)\,,~{}~{}~{}~{}m_{\chi}=938.513~{}\rm{MeV}\,,$ (34)
$Br(n\rightarrow\chi\phi)$ $\simeq$ $10^{8}~{}(\lambda_{\phi}\lambda_{q}\lambda_{\chi})^{2}\left(\frac{\rm{TeV}^{3}}{m_{\tilde{\chi}}M_{\Phi}^{2}}\right)\,,~{}~{}~{}~{}m_{\chi}=936.941~{}\rm{MeV}\,,$ (35)
and having $Br(n\rightarrow\chi\phi)\simeq 0.01$ would only require couplings $O(0.01-0.2)$.
(b) Decay into dark fermion and pair electron-positron: similarly, we can write the effective lagrangian:
${\cal L}_{n\rightarrow\chi e^{-}e^{+}}^{eff}=\kappa\bar{\chi}n\bar{e}e\,,$ (36)
with the 3-body decay rate given by:
$\Gamma_{n\rightarrow\chi e^{-}e^{+}}=\frac{\kappa^{2}m_{n}^{5}}{128\pi^{3}}\int_{4z^{2}}^{(1-x)^{2}}\frac{d\xi}{\sqrt{\xi}}(\xi-4z^{2})^{3/2}((1+x)^{2}-\xi)\sqrt{(1-x^{2}-\xi)^{2}-4\xi x^{2}}\,,$ (37)
where $x=m_{\chi}/m_{n}$ and $z=m_{e}/m_{n}$. The branching ratio of this channel has been constrained in Ref. Klopf:2019afh to be less than $10^{-4}$ for $e^{+}e^{-}$ kinetic energies in the range $37.5\,\rm{keV}\lesssim E_{e^{+}e^{-}}\lesssim 644\,\rm{keV}$ at 95% CL.
(c) Decay into dark fermion and electron-positron mediated by a light scalar: $n\rightarrow\chi+\phi\rightarrow\chi+e^{-}+e^{+}$
If we take $m_{\phi}\gtrsim 1~{}\rm{MeV}$, we re-introduce the decay channel into a pair electron-positron. The 3-body decay rate is given by:
$\frac{d\Gamma_{3}}{dE_{\chi}}=\frac{1}{64\pi^{3}m_{n}}|{\cal M}|^{2}\cdot\frac{1}{m_{23}^{2}}(E_{\chi}^{2}-m_{\chi}^{2})^{1/2}(m_{23}^{2}(m_{23}^{2}-4m_{e})^{2})^{1/2}\,,$ (38)
where $m_{23}^{2}=q^{2}=m_{n}^{2}-2E_{\chi}m_{n}+m_{\chi}^{2}$ is the momentum transferred, and ${\cal M}$ the scattering amplitude:
$|{\cal M}|^{2}=4\bar{\kappa}_{e}^{2}m_{n}(E_{\chi}+m_{\chi})(m_{n}^{2}+m_{\chi}^{2}-2m_{n}E_{\chi}-4m_{e}^{2})\cdot\frac{1}{(q^{2}-m_{\phi}^{2})^{2}+m_{\phi}^{2}\Gamma_{\phi}^{2}}\,.$ (39)
We have defined the coupling $\bar{\kappa}_{e}^{2}=\kappa_{\phi}^{2}h_{\phi}^{2}$, and included the Breit-Wigner propagator for the light scalar, with decay rate:
$\Gamma_{\phi}=\frac{h_{\phi}^{2}}{8\pi}\cdot m_{\phi}\cdot\left(1-\frac{4m_{e}^{2}}{m_{\phi}^{2}}\right)^{3/2}\,.$ (40)
Using the narrow width approximation,
$\frac{1}{(q^{2}-m_{\phi}^{2})^{2}+m_{\phi}^{2}\Gamma_{\phi}^{2}}\rightarrow\frac{\pi}{m_{\phi}\Gamma_{\phi}}\delta(q^{2}-m_{\phi}^{2})\,,$ (41)
we have for the 3-body decay rate:
[…]
with:
$\delta(m_{23}^{2}-m_{\phi}^{2})$ $=$ $\frac{1}{2m_{n}}\delta(\frac{m_{n}^{2}+m_{\chi}^{2}-m_{\phi}^{2}}{2m_{n}}-E_{\chi})\,,$ (43)
$F[E_{\chi},m_{n},m_{\chi}]$ $=$ $(E_{\chi}+m_{\chi})(m_{n}^{2}+m_{\chi}^{2}-2m_{n}E_{\chi}-4m_{e}^{2})\cdot(E_{\chi}^{2}-m_{\chi}^{2})^{1/2}\left(\frac{(m_{23}^{2}-4m_{e}^{2})}{m_{23}^{2}}\right)^{1/2}\,.$ (44)
Finally, we obtain:
$\Gamma_{3}=\frac{\bar{\kappa}_{e}^{2}}{16\pi^{2}}\cdot\frac{1}{m_{\phi}\Gamma_{\phi}}\cdot\frac{m_{\phi}^{2}}{2m_{n}}\cdot(1-\frac{4m_{e}^{2}}{m_{\phi}^{2}})^{3/2}\cdot(\tilde{E}_{\chi}-m_{\chi})^{1/2}(\tilde{E}_{\chi}+m_{\chi})^{3/2}\,,$ (45)
where we have defined: $\tilde{E}_{\chi}=(m_{n}^{2}+m_{\chi}^{2}-m_{\phi}^{2})/(2m_{n})$, $\tilde{E}_{\chi}+m_{\chi}=((m_{n}+m_{\chi})^{2}-m_{\phi}^{2})/(2m_{n})$, $\tilde{E}_{\chi}-m_{\chi}=((m_{n}-m_{\chi})^{2}-m_{\phi}^{2})/(2m_{n})$. And replacing the scalar decay rate Eq. (40) we have:
$\Gamma_{3}=\frac{\kappa_{\phi}^{2}}{16\pi}\cdot m_{n}\cdot\left(\frac{((m_{n}-m_{\chi})^{2}-m_{\phi}^{2})^{1/2}((m_{n}+m_{\chi})^{2}-m_{\phi}^{2})^{3/2}}{m_{n}^{2}}\right)\,,$ (46)
which is equal to the decay rate $\Gamma_{n\rightarrow\chi\phi}$ Eq. (32). If we want the model to explain the neutron decay anomaly, the branching ratio has to be O(1%), which requires either $m_{\phi}\lesssim 2m_{e}+0.0375$ MeV, or larger masses $m_{\phi}\gtrsim 2m_{e}+0.644$ MeV.